OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3? y = x + @bruhhh

OpenStudy (anonymous):

slope intercept is y-y1=m(x-x1) so y-2=3(x-6) :)

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

so you mean y=3x+6?

OpenStudy (anonymous):

Well if they want just the form of it, no simplifying or anything, then you'd leave it how I typed it. They just wanted you to plug the numbers in to form it

OpenStudy (anonymous):

it says y=___x+___

OpenStudy (anonymous):

so i can't put y-2

OpenStudy (anonymous):

Oh I gotchya, then they do want you to simplify it. y-2=3(x-6) y-2=3x-18 y=3x-16

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

they don't want me to simlify it

OpenStudy (anonymous):

i can't put y-2=3(x-6)

OpenStudy (anonymous):

That's really the only answer that would be right though if they have it set up as the normal equation of a line

OpenStudy (anonymous):

Oh I meant, put y=3x-16 instead lol

OpenStudy (anonymous):

I was just showing how they got from the point-slope form to the equation of a line, does that make sense at all?

OpenStudy (anonymous):

how you get 16 and the minus sign?

OpenStudy (anonymous):

So plugging in for point-slope gave us y-2=3(x-6) distributing the 3 to the x & 6 gives us y-2=3x-18 Then to get the y by itself, you add the 2 to both sides of the equal sign, giving us y=3x-16

OpenStudy (anonymous):

as i said it can't be y-2=3(x-6) because that's not the equation

OpenStudy (anonymous):

its y=___x+___

OpenStudy (anonymous):

Yes so it's now y=3x-16

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Lol it's the y-intercept can be negative if the value is negative. Hope that helped!!